OpenStudy (anonymous):

I really don't understand logarithms and my pre-cal final is on Thursday. Can someone help explain them to me?

OpenStudy (anonymous):

what don't you understand specifically ?

OpenStudy (anonymous):

it helps if you think of them as exponents. what kind of problems are you expected to do?

OpenStudy (anonymous):

solving trigonometric equations for one. I can do the simple problems, like rewriting in exponential form though I struggle. But I have a hard time with a long equation.

OpenStudy (anonymous):

Logs are the inverses of exponential functions in the form \[\log_{a}x \]

OpenStudy (anonymous):

confused. is it logs or trig?

OpenStudy (anonymous):

give us an example

OpenStudy (anonymous):

remember that \[b^x=y\iff \log_b(y)=x\] so for example \[2^x=10\iff \log_2(10)=x\]

OpenStudy (anonymous):

Oh, I'm so sorry. I mean logarithmic equations. I'm studying trigonometric identities too, I'm sorry. But an example would be log(base 8) (x^2 +11) = log(base 8) 92

OpenStudy (anonymous):

if the bases are the same, as they are in this question, then the inputs must be the same as well. so this really has very little to do with logs. solve \[x^2+11=91\]

OpenStudy (anonymous):

typo there i meant solve \[x^2+11=92\]

OpenStudy (anonymous):

so, you can just cancel the logs and just solve like a normal equation?

OpenStudy (anonymous):

As long as they are of the same base.

OpenStudy (anonymous):

yes but it your anser has to be >1

OpenStudy (anonymous):

it is not exactly "canceling" but logs are one to one functions, so if \[\log_b(x)=\log_b(y)\implies x = y\]

OpenStudy (anonymous):

I have yet to see one in my precal class where they aren't the same base. I don't think she would put that on the final, but who knows. If I come across a problem where they don't have the same base, what do I do?

OpenStudy (anonymous):

and as eherre said, make sure to only write the positive answer, because you cannot take the log of a negative number

OpenStudy (anonymous):

what about natural logs? For example, ln(3x^2-4) +ln(x^2 +1) = (2-x^2) ?

OpenStudy (anonymous):

if the bases are different you can use the "change of base formula" which says \[\log_a(x)=\frac{\log_b(x)}{\log_b(a)}\] which also means you can solve \[2^x=10\] by \[x=\frac{\log(10)}{\log(2)}\]

OpenStudy (anonymous):

okay, I understand you. What about condensing and expanding? I have done okay with the majority of those, but I have problems with the ones that have exponents.

OpenStudy (anonymous):

go to www.patrickjmt.com and he has a lot of helpful video tutorials